Solve the following equations
3x-4y=8
6x-2y=10
with x being 4/3 or -1
y being -1 or 4/3

Answer by Alan3354(69443)

Solve the following equations
3x-4y=8
6x-2y=10
I can do that.
--------------
with x being 4/3 or -1
y being -1 or 4/3
----------- I don't know what that means. x and y will be whatever they are.
3x-4y=8
6x-2y=10 --> 3x - y = 5
-----------
3x - 4y = 8
3x - y = 5
------------ Subtract
0x - 3y = 3
y = -1 yyyyyyyyyyyyyyyyyyyy
--------
3x -(-1) = 5
3x = 4
x = 4/3 xxxxxxxxxxxxxxxxxxx